OpenStudy (anonymous):

Find the sum. 1/2+3/8 A 1/2 B 3/4 C 7/8 D 4/10

OpenStudy (campbell_st):

use equivalent fractions and change 1/2 = 4/8 so its 4/8 + 3/8

OpenStudy (anonymous):

use equivalent fractions and change 1/2 = 4/8 so its 4/8 + 3/8 and 4/8 + 3/8 = 78

OpenStudy (anonymous):

7/8 is your answer